# Stoichiometry Representation Learning with Polymorphic Crystal Structures

### v1

@@ Abstract
Despite the recent success of machine learning (ML) in materials science, its success heavily relies on the structural description of crystal, which is itself computationally demanding and occasionally unattainable. Stoichiometry descriptors can be an alternative approach, which reveals the ratio between elements involved to form a certain compound without any structural information. However, it is not trivial to learn the representations of stoichiometry due to the nature of materials science called polymorphism, i.e., a single stoichiometry can exist in multiple structural forms due to the flexibility of atomic arrangements, inducing uncertainties in representation. To this end, we propose PolySRL, which learns the probabilistic representation of stoichiometry by utilizing the readily available structural information, whose uncertainty reveals the polymorphic structures of stoichiometry. Extensive experiments on sixteen datasets demonstrate the superiority of PolySRL, and analysis of uncertainties shed light on the applicability of PolySRL in real-world material discovery. The source code for PolySRL is available at https://github.com/Namkyeong/PolySRL_AI4Science.
@@ 1 Introduction
Recently, ML techniques have found their applications in the field of materials science to analyze the extensive amount of experimental and computational data available Zhang et al. (2023); Wang et al. (2023). However, the effectiveness of these ML models is not only influenced by the selection of appropriate models but also reliant on the numerical descriptors used to characterize the systems of interest. Although it is still an open problem to construct appropriate descriptions of materials, there is a general agreement on effective descriptors that encompass the following principles Huo & Rupp (2017); Faber et al. (2015); Bartók et al. (2013); Von Lilienfeld et al. (2015); Musil et al. (2021): Descriptors should 1) preserve the similarity or difference between two data points (preservativity), 2) be applicable to the entire materials domain of interest (versatility), and 3) be computationally more feasible to generate compared to computing the target property itself (computability).
Among various types of descriptors, there has been a notable surge of interest in using descriptors based on the knowledge of crystal structure in materials science. In particular, as shown in Figure 1(a), one can create graphical descriptions of crystalline systems by considering periodic boundary conditions and defining edges as connections between neighboring atoms within a specific distance Xie & Grossman (2018); Chen et al. (2019). However, these graphical descriptors depend on the structural details of crystals, which are usually obtained through computationally demanding and, in some cases, infeasible Density Functional Theory (DFT) calculations  Sholl & Steckel (2022). As a result, graphical descriptors are limited by the same computational bottleneck as DFT calculations, violating the principles of versatility and computability Damewood et al. (2023).
An alternative approach to using graphical descriptors is to develop material representations from stoichiometry [^1] Stoichiometry refers to the ratio between elements involved in a chemical reaction to form a compound. alone, as shown in Figure 1(b), which generates the representation of material solely based on its elemental composition Jha et al. (2018); Goodall & Lee (2020). Despite its simplicity, stoichiometry-based models have been shown to robustly offer a promising set of favorable elemental compositions for exploring new materials with cheap computational cost Damewood et al. (2023). However, this approach is inherently limited in that it overlooks the structural information of crystals, leading to inferior performance compared to graphical models Bartel et al. (2020) given that structural details strongly influence the crystal properties. This naturally prompts a question: “Is it possible for stoichiometry-based models to also capture the structural information of crystals?”
To answer the question, we propose a novel multi-modal representation learning framework for stoichiometry that incorporates readily available crystal structural information (i.e., stoichiometry and crystal structural information as multi-modal inputs), inspired by the recent success of multi-modal contrastive learning approaches in various domains Gan et al. (2022); Zong et al. (2023). For example, in computer vision, CLIP Radford et al. (2021) improves the zero-shot transferability of a vision model by matching captions and images. Moreover, 3D Infomax Stärk et al. (2022) improves 2D molecular graph representation in quantum chemistry by maximizing the mutual information with its corresponding 3D molecular representations.
However, naively adopting existing multi-modal contrastive learning approaches to the stoichiometry representation learning task is non-trivial due to the intrinsic characteristics of crystal structures, i.e., one-to-many relationship between stoichiometry and crystal structures stemming from the flexibility of atomic arrangements, which is also known as polymorphism. In other words, solely relying on stoichiometry would contradict the principle of preservativity, especially for polymorphic materials with the same stoichiometry. More specifically, polymorphism refers to the nature of a certain compound to exist in different crystallographic structures due to different arrangements of atoms, resulting in totally different physical, and chemical properties Bernstein (2020). An illustrative example of polymorphism is seen in the distinct forms of carbon: diamond and graphite (See Figure 1(c)). Diamond has a tetrahedral lattice structure with each carbon atom bonded to four others, resulting in its exceptional hardness and optical properties Che et al. (2000); Kidalov & Shakhov (2009). However, graphite has a planar layered structure where carbon atoms are bonded in hexagonal rings, forming sheets that can easily slide past each other, giving graphite its lubricating and conducting properties Wissler (2006); Jorio et al. (2008). Therefore, it is essential not only to obtain qualified stoichiometry representations, but also to account for the uncertainties stemming from polymorphism for real-world material discovery, which has been overlooked in previous studies Goodall & Lee (2020); Wang et al. (2021).
Figure 1: (a) Crystal structure of NaCl. (b), (c) Graphical and stoichiometry description of NaCl, respectively. (d) Diamond and Graphite share a single stoichiometry but have different structures.
@@
To this end, we propose Polymorphic Stoichiometry Representation Learning (PolySRL), which aims to learn the representation of stoichiometry as a probabilistic distribution of polymorphs instead of a single deterministic representation Oh et al. (2018); Chun et al. (2021). In particular, by assuming that polymorphs with an identical stoichiometry follow the same Gaussian distribution, PolySRL models each stoichiometry as a parameterized Gaussian distribution with learnable mean and variance vectors, whose distribution is trained to cover the range of polymorphic structures in representation space. By doing so, we expect the mean of Gaussian distribution serves as the representation of the stoichiometry, and the variance reflects the uncertainty stemming from the existence of various polymorphic structures, enabling PolySRL to assess the degree to which the representation adheres to the principle of preservativity. In this work, we make the following contributions:
@@
  • Recognizing the advantages and limitations of both structural and stoichiometry descriptors, we propose a multi-modal representation learning framework for stoichiometry, called PolySRL, which incorporates structural information of crystals into stoichiometry representations.
@@
  • To capture uncertainties of stoichiometry stemming from various polymorphs, PolySRL learns a probabilistic representation for each stoichiometry instead of a deterministic representation.
@@
  • Extensive experiments on sixteen datasets demonstrate the superiority of PolySRL in learning representation of stoichiometry and predicting its physical properties. Moreover, we observe that measured uncertainties reflect various challenges in materials science, highlighting the applicability of PolySRL for real-world material discovery.
@@
To the best of our knowledge, this is the first work that learns generalized representations of stoichiometry by simultaneously considering the crystal structural information and the polymorphism as uncertainty, which is crucial for the process of real-world material discovery.
2 Related Works
2.1 Graph Neural Networks for Materials
Among various ML methods, graph neural networks (GNNs) have been rapidly adopted by modeling crystal structures as graphical descriptions inspired by the recent success of GNNs in biochemistry Gilmer et al. (2017); Stokes et al. (2020); Jiang et al. (2021); Huang et al. (2022); Lee et al. (2023b, a). Specifically, CGCNN Xie & Grossman (2018) first proposes a message-passing framework based on a multi-edge graph to capture interactions across cell boundaries, resulting in highly accurate prediction for eight distinct material properties. Building upon this multi-edge graph foundation, MEGNet Chen et al. (2019) predicts various crystal properties by incorporating a physically intuitive strategy to unify multiple GNN models. Moreover, ALIGNN Choudhary & DeCost (2021) proposes to utilize a line graph, in addition to a multi-edge graph, to model additional structural features such as bond angles and local geometric distortions. Despite the recent success of graph-based approaches, their major restriction is the requirement of atomic positions, which are typically determined through computationally intensive and sometimes infeasible DFT calculations. As a result, their effectiveness is mainly demonstrated in predicting properties for systems that have already undergone significant computational effort, restricting their utility in the materials discovery workflow Damewood et al. (2023).
2.2 Stoichiometry Representation Learning
Material representations can be alternatively constructed solely based on stoichiometry, which indicates the concentration of the constituent elements, without any knowledge of the crystal structure Damewood et al. (2023). While stoichiometry has historically played a role in effective materials design Callister & Rethwisch (1964); Pauling (1929), it has been recently demonstrated that deep neural networks (DNNs) tend to outperform conventional approaches when large datasets are available. Specifically, ElemNet Jha et al. (2018) takes elemental compositions as inputs and trains DNNs with extensive high-throughput OQMD dataset Kirklin et al. (2015), showing improvements in performance as the network depth increases, up to a point where it reaches 17 layers. Roost Goodall & Lee (2020) utilizes GNNs for stoichiometry representation learning by creating a fully connected graph in which nodes represent elements, allowing for the modeling of interactions between these elements. Instead of the message-passing scheme, CrabNet Wang et al. (2021) introduces a self-attention mechanism to adaptively learn the representation of individual elements based on their chemical environment. While these methods are trained for a specific task, PolySRL aims to learn generalized stoichiometry representations for various tasks considering 1) the structural information and 2) polymorphism in crystal, both of which have not been explored before.
2.3 Probabilistic Representation Learning
First appearing in 2014 with the introduction of probabilistic word embeddings Vilnis & McCallum (2014), probabilistic representations got a surge of interest from ML researchers by offering numerous benefits in modeling uncertainty pertaining to a representation. Specifically, in the computer vision domain, Shi & Jain (2019) proposes to probabilistically represent face images to address feature ambiguity in real-world face recognition. Moreover, Oh et al. (2018) introduces Hedged Instance Embeddings (HIB), which computes a match probability between point estimates but integrates it over the predicted distributions via Monte Carlo estimation. This idea has been successfully extended to cross-modal retrieval Chun et al. (2021), video representation learning Park et al. (2022), and concept prediction Kim et al. (2023). In this paper, we aim to learn a probabilistic representation of stoichiometry, where the uncertainties account for various polymorphs associated with a single stoichiometry, enhancing the reliability of the material discovery process.
3 Preliminaries
3.1 Stoichiometry Graph Construction
@@
Given a stoichiometry, we use =$\mathcal{E}=\{e_{1},\ldots,e_{n_{e}}\}$ to denote its unique set of elements, and =$\mathcal{R}=\{r_{1},\ldots,r_{n_{e}}\}$ to denote the compositional ratio of each element in the stoichiometry. We construct a fully connected stoichiometry graph $\mathcal{G}^{a}=(\mathcal{E},\mathcal{R},\mathbf{A}^{a})$, where $\mathbf{A}^{a}\in\{1\}^{{n_{e}}\times{n_{e}}}$ indicates the adjacency matrix of a fully connected graph Goodall & Lee (2020). Then, we adopt GNNs as the stoichiometry encoder $f^{a}$, which aims to learn the stoichiometry representation by capturing complex relationships between elements via the message-passing scheme. Additionally, $\mathcal{G}^{a}$ is associated with an elemental feature matrix $\mathbf{X}^{a}\in\mathbb{R}^{{n_{e}}\times F}$ where $F$ is the number of features.
3.2 Structural Graph Construction
@@
Given a crystal structure $(\mathbf{P},\mathbf{L})$, suppose the unit cell has $n_{s}$ atoms, we have $\mathbf{P}={[\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{n_{s}}]}^{\intercal}\in\mathbb{R}^{{n_{s}}\times 3}$ indicating the atom position matrix and $\mathbf{L}={[\mathbf{l}_{1},\mathbf{l}_{2},\mathbf{l}_{3}]}^{\intercal}\in\mathbb{R}^{3\times 3}$ representing the lattice parameter describing how a unit cell repeats itself in three directions. Based on the crystal parameters, we construct a multi-edge graph $\mathcal{G}^{b}=(\mathcal{V},\mathbf{A}^{b})$ that captures atom interactions across cell boundaries Xie & Grossman (2018). Specifically, $v_{i}\in\mathcal{V}$ denotes an atom $i$ and all its duplicates in the infinite 3D space whose positions are included in the set $\{\hat{\mathbf{p}}_{i}|\hat{\mathbf{p}}_{i}=\mathbf{p}_{i}+k_{1}\mathbf{l}_{1}+k_{2}\mathbf{l}_{2}+k_{3}\mathbf{l}_{3},k_{1},k_{2},k_{3}\in\mathbb{Z}\}$, where $\mathbb{Z}$ denotes the set of all the integers. Moreover, $\mathbf{A}^{b}\in\{0,1\}^{n_{s}\times n_{s}}$ denotes an adjacency matrix, where $\mathbf{A}_{i,j}^{b}=1$ if two atoms $i$ and $j$ are within the predefined radius $r$ and $\mathbf{A}^{b}_{ij}=0$ otherwise. Moreover, a single stoichiometry graph $\mathcal{G}^{a}$ is associated with a set of polymorphic crystal structural graphs $\mathcal{P}^{\mathcal{G}^{a}}$, i.e., $\mathcal{P}^{\mathcal{G}^{a}}=\{\mathcal{G}_{1}^{b},\ldots,\mathcal{G}_{n_{p}}^{b}\}$, where $n_{p}$ is the number of polymorphs for the stoichiometry. Note that each node in $\mathcal{G}^{b}$ is associated with a learnable feature $\mathbf{x}^{b}\in\mathbb{R}^{F}$, which is shared across all crystals, to make sure we utilize only structural information. We provide further details on structural graph construction in Appendix A.
3.3 Task Descriptions
@@
Given the stoichiometry graph $\mathcal{G}^{a}$ and the structural graph $\mathcal{G}^{b}$ of a single crystal, our objective is to acquire a stoichiometry encoder denoted as $f^{a}$, alongside mean and variance modules referred to as $f_{\mu}^{a}$ and $f_{\sigma}^{a}$, which associate structural information of $\mathcal{G}^{b}$ into latent representation of stoichiometry $\mathcal{G}^{a}$. Then, the modules are applied to a range of downstream tasks, a scenario frequently encountered in real-world material science where solely stoichiometry of material is accessible.
4 Methodology: PolySRL
In this section, we present Polymorphic Stoichiometry Representation Learning (PolySRL), which learns the representation of stoichiometry regarding polymorphic structures of crystals. Overall model architecture is illustrated in Figure 2.
4.1 Structural Graph Encoder
@@
While structural information plays an important role in determining various properties of crystals, previous studies have overlooked the readily available crystal structures Jain et al. (2013) for stoichiometry representation learning Jha et al. (2018); Goodall & Lee (2020); Wang et al. (2021). To this end, we use a GNN encoder to learn the representation of crystal structure, which is expected to provide guidance for learning the representation of stoichiometry. More formally, given the crystal structural graph $\mathcal{G}^{b}=(\mathbf{x}^{b},\mathbf{A}^{b})$, we obtain a structural representation of a crystal as follows:
$\mathbf{z}^{b}=\text{Pooling}(\mathbf{Z}^{b}),\,\,\,\,\mathbf{Z}^{b}=f^{b}(\mathbf{x}^{b},\mathbf{A}^{b}),$ (1)
@@
where $\mathbf{Z}^{b}\in\mathbb{R}^{n_{s}\times F}$ is a matrix whose each row indicates the representation of each atom in the crystal structure, $\mathbf{z}^{b}$ indicates the latent representation of a crystal structure, and $f^{b}$ is the GNN-based crystal structural encoder. In this paper, we adopt graph networks Battaglia et al. (2018) as the encoder, which is a generalized version of various GNNs, and sum pooling is used as the pooling function. We provide further details on the GNNs in Appendix B.1.
Figure 2: Overall model architecture.
4.2 Probabilistic Stoichiometry Encoder

### v2

# Compositional Representation of Polymorphic Crystalline Materials

@@ Abstract
Machine learning (ML) has seen promising developments in materials science, yet its efficacy largely depends on detailed crystal structural data, which are often complex and hard to obtain, limiting their applicability in real-world material synthesis processes. An alternative, using compositional descriptors, offers a simpler approach by indicating the elemental ratios of compounds without detailed structural insights. However, accurately representing materials solely with compositional descriptors presents challenges due to polymorphism, where a single composition can correspond to various structural arrangements, creating ambiguities in its representation. To this end, we introduce PCRL, a novel approach that employs probabilistic modeling of composition to capture the diverse polymorphs from available structural information. Extensive evaluations on sixteen datasets demonstrate the effectiveness of PCRL in learning compositional representation, and our analysis highlights its potential applicability of PCRL in material discovery. The source code for PCRL is available at https://github.com/Namkyeong/PCRL.
@@ 1 Introduction
Figure 1: (a) A crystal can have multiple descriptors for ML input, such as graphical and compositional descriptors. (b) Diamond and Graphite are polymorphic crystal structures of composition C. While these crystal structures share the same composition, their properties are completely different. (c) Graphical descriptors necessitate costly recalculation of crystal structures during the material synthesis process, thereby restricting the ML capability to the same bottleneck as conventional materials discovery processes. (d) Compositional representation enables the use of ML without requiring these expensive structural calculations. (e) Pre-training composition graph encoder encoder with contrastive learning. While the structural graph encoder obtains a deterministic structural representation of crystal, the probabilistic composition encoder learns to represent each composition as a parameterized probabilistic distribution by acquiring mean and diagonal covariance matrices. Both encoders are jointly trained with soft contrastive loss in representation space. (f) The pre-trained composition mean encoder $f_{\mu}^{a}$ can be utilized to predict various properties of materials, while the composition uncertainty encoder $f_{\sigma}^{a}$ can guide scientists in determining which materials to investigate further.
Recently, machine learning (ML) has increasingly been leveraged in materials science to sift through the abundant experimental and computational data available [1, 2]. Interest has particularly spiked in graphical descriptors derived from the structural properties of crystals. Notably, graphical representations can be created by incorporating periodic boundary conditions and connecting neighboring atoms within a certain range, as depicted in Figure 1 (a) [3, 4]. However, in the real-world material synthesis process, the reorganization of atomic arrangements frequently occurs due to the mixing of ingredients, heat treatments, and chemical reactions with solvents, necessitating the recalculation of crystal structures. These recalculations rely on the computationally intensive and sometimes infeasible Density Functional Theory (DFT) [5], limiting the ML capability to the same computational bottleneck as DFT calculations. In other words, while ML has garnered significant interest from material scientists because of its efficiency and speed compared to DFT calculations, it still necessitates DFT calculations when utilizing graphical representations of materials in ML models during the materials discovery process. Moreover, in materials science, doping and alloying are used to improve the performance of materials such as semiconductors  [6, 7]. However, because doping and alloying chaotically change the structure of pure materials, it can lead to cases where calculations of the structure are infeasible, limiting the use of graphical representation for materials.
An alternative to graphical descriptors is to utilize material representations based purely on composition, which denotes the elemental ratios present in a chemical compound, as illustrated in Figure 1 (a) [8, 9]. While straightforward, models based on composition have proven to be effective, offering a promising array of potential elemental combinations for new material discovery at a low computational expense[10]. Nonetheless, this method has a significant drawback as it ignores the structural aspects of crystals, which can result in less effective performance compared to graphical models [11]. This discrepancy emphasizes the structural impact on crystal properties and raises an important inquiry: “Can composition-based models be enhanced by incorporating structural details of crystals?
Recent multi-modal contrastive pre-training strategies in various domains may provide an answer to this question [12, 13]. Specifically, the CLIP model [14] in computer vision significantly enhances zero-shot learning by aligning images with their captions. Moreover, 3D Infomax [15] improves the accuracy of 2D molecular graphs in quantum chemistry by maximizing their mutual information with 3D molecules during the pre-training step and then using only 2D molecular graphs during the fine-tuning step. Similarly, one could introduce both compositional and structural information as multiple modalities for materials during the pre-training stage, and then use only the compositional information during the fine-tuning stage.
However, implementing multi-modal contrastive learning techniques directly for compositional representation learning encounters obstacles. This is due to the inherent nature of crystal structures, named polymorphism, where a single composition can exist in multiple crystal structures. That is, the relationship between composition and crystal structures is one-to-many rather than one-to-one, leading to intrinsic uncertainties in practical applications.
More specifically, polymorphism describes how a single compound can form various crystallographic structures due to different atomic arrangements, which leads to vastly different physical and chemical properties [16]. For example, carbon can exist as either diamond or graphite as shown in Figure 1 (b). In diamond, carbon atoms are arranged in a tetrahedral lattice with each carbon atom bonded to four others, giving it extraordinary hardness and unique optical properties [17, 18]. In contrast, graphite has a planar layered structure where carbon atoms are bonded in hexagonal rings, allowing the sheets to slide over each other easily, which results in its lubricating and electrical conductivity properties [19, 20]. Although the compositional models give up the ability to handle such polymorphs while circumventing the structure bottleneck [9], one can alternatively provide uncertainties caused by polymorphism, which is crucial in real-world materials discovery.
1.1 Our Approach
@@
To this end, we propose a multi-modal pre-training framework for crystalline materials, namely Probabilistic Compositional Representation Learning (PCRL). This framework aims to pre-train a composition encoder that can be universally applied to various downstream tasks. Specifically, the composition encoder outputs representation of composition as a probabilistic distribution of polymorphs instead of a single deterministic representation [21, 22]. In particular, by assuming that polymorphs with an identical composition follow the same Gaussian distribution, PCRL models each composition as a parameterized Gaussian distribution with learnable mean and variance vectors, whose distribution is trained to cover the range of polymorphic structures in representation space. By doing so, we expect the mean of the parameterized Gaussian distribution to represent the composition and the variance to reflect the uncertainty stemming from the existence of various polymorphic structures.
In this work, we make the following contributions:
@@
  • Recognizing the advantages and limitations of both structural and compositional descriptors, we propose a multi-modal representation learning framework for composition, called PCRL, which incorporates structural information of crystals into compositional representations.
  • Unlike previous material property prediction works that rely on composition, our goal is to pre-train a universally applicable composition encoder that can be used for predicting various properties.
@@
  • To capture uncertainties of composition stemming from various polymorphs, PCRL learns a probabilistic representation for each composition, whose mean serves as the representation of the composition and variance reflects uncertainty stemming from the polymorphism.
@@
  • Extensive experiments on sixteen datasets demonstrate the superiority of PCRL in learning universally applicable representation of composition and predicting its various physical properties. Moreover, measured uncertainties reflect various challenges in materials science, highlighting the applicability of PCRL for real-world materials discovery.
@@
To the best of our knowledge, this is the first work that learns universal compositional representations by simultaneously considering the crystal structural information and the polymorphism as uncertainty, which is crucial for real-world material discovery.
2 Method
2.1 Preliminaries
2.1.1 Compositional Graph Construction
@@
Given a composition, we use =$\mathcal{E}=\{e_{1},\ldots,e_{n_{e}}\}$ to denote its unique set of elements, and =$\mathcal{R}=\{r_{1},\ldots,r_{n_{e}}\}$ to denote the compositional ratio of each element in the composition. We construct a fully connected compositional graph $\mathcal{G}^{a}=(\mathcal{E},\mathcal{R},\mathbf{A}^{a})$, where $\mathbf{A}^{a}\in\{1\}^{{n_{e}}\times{n_{e}}}$ indicates the adjacency matrix of a fully connected graph [9]. Then, we adopt GNNs as the composition encoder $f^{a}$, which aims to learn the compositional representation by capturing complex relationships between elements via the message-passing scheme. Additionally, $\mathcal{G}^{a}$ is associated with an elemental feature matrix $\mathbf{X}^{a}\in\mathbb{R}^{{n_{e}}\times F}$ where $F$ is the number of features.
2.1.2 Structural Graph Construction
@@
Given a crystal structure $(\mathbf{P},\mathbf{L})$, suppose the unit cell has $n_{s}$ atoms, we have $\mathbf{P}={[\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{n_{s}}]}^{\intercal}\in\mathbb{R}^{{n_{s}}\times 3}$ indicating the atom position matrix and $\mathbf{L}={[\mathbf{l}_{1},\mathbf{l}_{2},\mathbf{l}_{3}]}^{\intercal}\in\mathbb{R}^{3\times 3}$ representing the lattice parameter describing how a unit cell repeats itself in three directions. Based on the crystal parameters, we construct a multi-edge graph $\mathcal{G}^{b}=(\mathcal{V},\mathbf{A}^{b})$ that captures atom interactions across cell boundaries [3]. Specifically, $v_{i}\in\mathcal{V}$ denotes an atom $i$ and all its duplicates in the infinite 3D space whose positions are included in the set $\{\hat{\mathbf{p}}_{i}|\hat{\mathbf{p}}_{i}=\mathbf{p}_{i}+k_{1}\mathbf{l}_{1}+k_{2}\mathbf{l}_{2}+k_{3}\mathbf{l}_{3},k_{1},k_{2},k_{3}\in\mathbb{Z}\}$, where $\mathbb{Z}$ denotes the set of all the integers. Moreover, $\mathbf{A}^{b}\in\{0,1\}^{n_{s}\times n_{s}}$ denotes an adjacency matrix, where $\mathbf{A}_{i,j}^{b}=1$ if two atoms $i$ and $j$ are within the predefined radius $r$ and $\mathbf{A}^{b}_{ij}=0$ otherwise. Furthermore, a single compositional graph $\mathcal{G}^{a}$ is associated with a set of polymorphic crystal structural graphs $\mathcal{P}^{\mathcal{G}^{a}}$, i.e., $\mathcal{P}^{\mathcal{G}^{a}}=\{\mathcal{G}_{1}^{b},\ldots,\mathcal{G}_{n_{p}}^{b}\}$, where $n_{p}$ is the number of polymorphs for the composition. We provide further details on structural graph construction in Appendix A.
2.1.3 Task Description
@@
Given the compositional graph $\mathcal{G}^{a}$ and the structural graph $\mathcal{G}^{b}$ of a single crystal, our objective is to acquire a composition encoder denoted as $f^{a}$, alongside mean and variance modules referred to as $f_{\mu}^{a}$ and $f_{\sigma}^{a}$. Then, these modules including pre-trained composition encoder $f^{a}$, and the mean $f_{\mu}^{a}$ and variance $f_{\sigma}^{a}$ modules, are employed in a range of downstream tasks, a scenario frequently encountered in real-world material synthesizing processes where solely composition of crystalline material is accessible.
2.2 Structural Graph Encoder $f^{b}$
@@
While structural information plays an important role in determining various properties of crystals, previous studies have overlooked the readily available crystal structures [23] for compositional representation learning [8, 9, 24]. To this end, we use a GNN encoder to learn the representation of crystal structure, which is expected to provide guidance for learning the representation of composition. More formally, given the crystal structural graph $\mathcal{G}^{b}=(\mathbf{x}^{b},\mathbf{A}^{b})$, we obtain a structural representation of a crystal as follows:
$\small\mathbf{z}^{b}=\text{Pooling}(\mathbf{Z}^{b}),\,\,\,\,\mathbf{Z}^{b}=f^{b}(\mathbf{x}^{b},\mathbf{A}^{b}),$ (1)
@@
where $\mathbf{Z}^{b}\in\mathbb{R}^{n_{s}\times F}$ is a matrix whose each row indicates the representation of each atom in the crystal structure, $\mathbf{z}^{b}$ indicates the latent representation of a crystal structure, and $f^{b}$ is the GNN-based crystal structural encoder. In this paper, we adopt graph neural networks [25] as the encoder, which is a generalized version of various GNNs, and sum pooling is used as the pooling function. We provide further details on the GNNs in Appendix B.1.
2.3 Probabilistic Composition Encoder $f^{a}$